OpenStudy (anonymous):

A firecracker shoots up from a hill 160 feet high with an initial speed of 90 feet per second. Using formula H(t)= -16t^2 + vt + s approximately how long will it take the firecracker to hit the ground?

OpenStudy (anonymous):

solve for \(t\) \[-16t^2+90t+160=0\]

OpenStudy (jhannybean):

factor out a 16.

OpenStudy (anonymous):

Jhannybean I have and my answer is not correct. Satellite: I have redone that 3x and my answer is not from the chosen answers.

OpenStudy (anonymous):

anymore help??

sam (.sam.):

s will be the height as usual, v is the velocity H(t)= 0 when it touches the ground H(t)=-16t^2 + vt + s=0 Then \[-16t^2+90t+160=0\] Solve for t

sam (.sam.):

Did you use quadratic formula?

sam (.sam.):

\[t=\frac{-90 \pm \sqrt{90^2-4(-16)(160)}}{2(-16)}\]

OpenStudy (anonymous):

yes. I have had to move on. I cant get the correct answer. Thank you so much.

OpenStudy (jhannybean):

You would have had two answers for t, one solving for [-b + sqrt{b^2-4(a)(c)}]/2a and the other one for [-b-sqrt{b^2-4ac}]/2a